# Channel-wise Parallelizable Spiking Neuron with Multiplication-free Dynamics and Large Temporal Receptive Fields

### v1

@@ Abstract
Spiking Neural Networks (SNNs) are distinguished from Artificial Neural Networks (ANNs) for their sophisticated neuronal dynamics and sparse binary activations (spikes) inspired by the biological neural system. Traditional neuron models use iterative step-by-step dynamics, resulting in serial computation and slow training speed of SNNs. Recently, parallelizable spiking neuron models have been proposed to fully utilize the massive parallel computing ability of graphics processing units to accelerate the training of SNNs. However, existing parallelizable spiking neuron models involve dense floating operations and can only achieve high long-term dependencies learning ability with a large order at the cost of huge computational and memory costs. To solve the dilemma of performance and costs, we propose the mul-free channel-wise Parallel Spiking Neuron, which is hardware-friendly and suitable for SNNs’ resource-restricted application scenarios. The proposed neuron imports the channel-wise convolution to enhance the learning ability, induces the sawtooth dilations to reduce the neuron order, and employs the bit shift operation to avoid multiplications. The algorithm for design and implementation of acceleration methods is discussed meticulously. Our methods are validated in neuromorphic Spiking Heidelberg Digits voices, sequential CIFAR images, and neuromorphic DVS-Lip vision datasets, achieving the best accuracy among SNNs. Training speed results demonstrate the effectiveness of our acceleration methods, providing a practical reference for future research.
111$\dagger$ Corresponding authors
@@ 1 Introduction
Inspired by the biological neural system, Spiking Neural Networks (SNNs) are regarded as the third generation neural network models (Maass, 1997). By emulating neuronal dynamics and spike-based communication characteristics in the brain (Tavanaei et al., 2019), SNNs effectively capture temporal information and achieve event-driven efficient computation, providing a novel paradigm for building the spike-based machine intelligence (Yao et al., 2024b).
@@
Spiking neurons are the key component that distinguishes SNNs from Artificial Neural Networks (ANNs) (Li et al., 2024a). They integrate input currents from synapses to membrane potentials by complex neuronal dynamics and fire spikes when the membrane potentials cross the threshold. These proceedings are generally described by several discrete-time difference equations (Fang et al., 2021b, 2023a) in a formulation similar to the Recurrent Neural Networks. The discrete threshold-triggered firing mechanism induces the nondifferentiable problem and impedes the application of gradient descent methods. Recently, this problem has been resolved to a considerable degree by the surrogate gradient methods (Wu et al., 2018; Shrestha and Orchard, 2018; Neftci et al., 2019).
@@
Deep SNNs (Fang et al., 2021a; Zhou et al., 2023; Yao et al., 2024a) commonly use stateless synapses, i.e. the weights are shared across time-steps and outputs only depend on the inputs at the same time-step, to extract spatial features. Consequently, dynamic spiking neurons play the critical role in SNNs to extract temporal features and have attracted many research interests. Most of the previous research focuses on increasing the neuron model complexity with learnable parameters (Fang et al., 2021b; Yao et al., 2022) or adaptive dynamics (Huang et al., 2024a), which strengths the model capacity but brings extra computation costs. Another issue is that traditional spiking neuron models run in a serial step-by-step mode and cannot fully utilize the powerful massive parallel computing ability of Graphics Processing Units (GPUs), which leads to slower training speed of SNNs than ANNs. Recently, parallelizable spiking neuron models (Fang et al., 2023b; Yarga and Wood, 2023; Huang et al., 2024c; Chen et al., 2024a) have been proposed that overwhelm traditional serial models in running speed, showing a promising solution to accelerate the training of SNNs.
@@
In previous designs of spiking neurons, the computational cost associated with the neuron dynamics is generally overlooked. For instance, the Complementary Leaky Integrated-and-Fire (CLIF) neuron (Huang et al., 2024a) introduces the computationally expensive sigmoid exponentiation, the Parallel Spiking Neuron (PSN), and the masked PSN (Fang et al., 2023b) rely on the dense floating-point matrix multiplication. Compared to PSN and masked PSN, sliding PSN (Fang et al., 2023b) only requires convolutional operations and demonstrates superior performance in handling time series with variable lengths. However, according to the study by (Fang et al., 2023b), sliding PSN only achieves comparable performance to PSN when using a large convolutional kernel size, denoted as the order of the neuron, which significantly increases computational cost and memory usage.
@@
  • To enhance the temporal information capturing ability, we derive the sliding PSN by applying the channel-wise separable convolutions. To solve the dilemma of large temporal receptive fields and computational costs, we use the dilated convolution. Compared to sliding PSN, our improvement does not introduce any additional floating point operations (FLOPs).
  • To avoid the costly multiplication operations, we supersede them with bit-shift operations, which are hardware-friendly for resource-restricted neuromorphic chips. The theoretical energy cost and area for hardware implementation with 8-bit integers (INT8) precision under 45nm CMOS is reduced 8$\times$, from 0.2 $pJ$ and 282 $\mu m^{2}$ to 0.024 $pJ$ and 34 $\mu m^{2}$ (You et al., 2024), respectively.
  • We deliberate on the implementations of SNNs with mul-free channel-wise PSN on GPUs for efficient training. We propose an autoselect algorithm to choose the fastest implementations, which is practical for future research about accelerating parallelizable spiking neurons.
  • We achieve SOTA performance on various temporal tasks, validating the superior capability of the proposed methods in learning long-term dependencies.
@@ 2 Related Work
2.1 Hardware-friendly Network Design
@@
To deploy neural network models to edge devices such as mobile phones and Field Programmable Gate Arrays with limited energy, memory, and computational ability, a promising solution is hardware-friendly network design. Various methodologies have been proposed. Network quantization (Gholami et al., 2021) quantizes the original weights and activations to low bits. Typical models in 8-bit integers require 4×4$\times$4 × less memory consumption and achieve up to 4×4$\times$4 × faster computation than models in 32-bit floating-points. Network pruning (Cheng et al., 2024) prunes synapses and neurons to reduce the size of the model. Classic pruning methods can achieve 4×4$\times$4 × compression ratio for ResNet-18 on ImageNet with about $5\%$ accuracy drop (Blalock et al., 2020). Knowledge distillation (Gou et al., 2021) employs a large teacher network to supervise the learning of a small student network, and the student network can obtain a competitive or even higher performance than the teacher network. Apart from the above universal methods, SNNs (Maass, 1997; Roy et al., 2019) achieve extreme power efficiency by asynchronous event-driven computation in tailored neuromorphic chips. For instance, Intel Loihi (Davies et al., 2018) consumes 48×48$\times$48 × energy efficiency than CPUs in solving the LASSO optimization problem; Tsinghua Tianjic (Pei et al., 2019) achieves up to $10^{4}$ times power efficiency over the Titan-Xp GPU when classifying the NMNIST (Orchard et al., 2015).
2.2 Spiking Deep Learning
The performance of SNNs surges by the induction of deep learning methods. ANN to SNN conversion (Cao et al., 2015; Hu et al., 2023) and surrogate gradient learning (Neftci et al., 2019) are two primary methods for spiking deep learning (Chollet, 2017). The conversion method leverages the average firing rate of spiking neurons to approximate the continuous activations in ANNs, transforming a pre-trained ANN into an SNN through techniques including threshold adjustment and weight normalization. However, the conversion method requires a substantial number of time-steps to estimate accurate firing rates. The surrogate gradient learning method employs the derivative of a continuous function to replace the derivative of the Heaviside function used in the spike generation process. This approach enables SNNs to achieve credit assignment through backpropagation through time (BPTT) and gradient descent. Compared with the conversion method, the surrogate gradient method requires much fewer time-steps, but the training cost is higher due to the use of BPTT.
2.3 Spiking Neuron Models
@@
The improvement of spiking neuron models provides a general method to upgrade SNNs, which attracts many interests in the research community. The Parametric Leaky Integrated-and-Fire (PLIF) spiking neuron (Fang et al., 2021b) parameterizes the membrane time constant $\tau_{m}$ by a sigmoid function and can learn $\tau_{m}$ by gradient descent during training, showing better accuracy and lower latency than the traditional Leaky Integrated-and-Fire (LIF) neuron with fixed $\tau_{m}$. The Gated LIF (GLIF) neuron (Yao et al., 2022) assembles the learnable gate units to fuse different bio-features in the neuronal behaviors of membrane leakage, integration accumulation, and reset, achieving impressive performance by these rich neuronal patterns. The Complementary LIF (CLIF) neuron (Huang et al., 2024b) introduces the complementary membrane potential into the LIF neuron, effectively capturing and maintaining information related to membrane potential decay. However, the sigmoid used in its neuronal dynamics cannot be removed during inference, which is costly for neuromorphic chips.
The Parallel Spiking Neuron (PSN) family (Fang et al., 2023b) and the Stochastic Parallelizable Spiking Neuron (Stochastic PSN) (Yarga and Wood, 2023) are the first parallelizable spiking neuron models. These two models convert the iterative neuronal dynamics to a non-iterative formulation by removing the neuronal reset. Extending from PSN, several variants are proposed. The Parallel Multi-compartment Spiking Neuron (PMSN) (Chen et al., 2024b) introduces multiple interacting substructures to enhance the learning ability over diverse timescales. The Parallel Spiking Unit (PSU) (Li et al., 2024b) adds a fully-connected layer with sigmoid activations inside the neuron to approximate the neuronal reset. These methods obtain performance gains over PSN in certain datasets, but increase the complexity of neuron models and slow down training speeds.
@@ 3 Preliminary
3.1 Vanilla Spiking Neuron
The general formulations of spiking neurons are as follows:
@@ 3.1 Vanilla Spiking Neuron
$V[t]$ =$=\begin{cases}H[t]\cdot(1-S[t])+V_{reset}\cdot S[t],\text{hard reset}\\ H[t]-V_{th}\cdot S[t],\text{soft reset}\end{cases}.$ (3)
Eq.(1) illustrates the neuronal charging process, where $X[t]$ is the input current, $H[t],V[t]$ are the membrane potential before charging and after resetting, and $f$ is the charging equation specified for different spiking neurons. After charging, the membrane potential $H[t]$ will be compared with the threshold $V_{th}$, described as Eq.(2) to determine whether firing. $\Theta(x)$ is the Heaviside step function, defined as $\Theta(x)=1$ for $x\geq 0$ and $\Theta(x)=0$ for $x<0$. If firing, the membrane potential will be reset as in Eq.(3). There are mainly two types of reset methods: hard reset will force the $V[t]$ to $V_{reset}$, while soft reset will subtract $V_{th}$ from $V[t]$.
@@ 3.2 Parallel Spiking Neuron
Fang et al. (2023b) found that for commonly used spiking neurons with a linear sub-threshold dynamic Eq.(1), such as the Integrate-and-Fire (IF) neuron and the LIF neuron, the neuronal dynamics could be expressed in a non-iterative form after removing the reset equation Eq.(3):
@@
Fang et al. (2023b) extended Eq.(4) by setting $W[t][i]$ as a learnable parameter, and proposed the PSN with following neuronal dynamics:
@@
where $T$ is the sequence length. For simplicity, we ignore the batch dimension. The PSN does not involve iteration over time-steps. The core computation of the PSN is the matrix multiplication, which is highly optimized on GPUs and can be computed in parallel. Modified from the PSN, the sliding PSN is proposed by (Fang et al., 2023b) with hidden states generating from the last $k$ inputs by a shared weight $\bm{W}\in\mathbb{R}^{k}$ across time-steps, whose neuronal dynamics are as follows:
@@ 4.1 Channel-wise and Dilated Convolution
In PSN and sliding PSN, the weights of the neurons are shared across all channels. However, the visualization of feature maps from an SNN conducted by (Fang et al., 2021b) implies that the difference between channels is huge, e.g., one channel extracts the edges and another channel extracts the background at all time-steps (refer to Figure S4 in (Fang et al., 2021b) for more details). This coarse design concept may fail to capture the subtle disparity of features in channels. To solve this issue, we extend the weights to channel-wise.
To capture long-term dependency, the sliding PSN must use a large order $k$, resulting in a significant rise in computational cost and memory usage. We overcome this issue through the dilated convolution (Yu and Koltun, 2016), where the convolution no longer processes consecutive inputs $(...,X[t-2],X[t-1],X[t])$, but instead $(...,X[t-2d],X[t-d],X[t])$, with $d>1$ as the dilation rate.
@@
where $\bm{W}\in\mathbb{R}^{C\times k}$ is the learnable weight, $k$ is the order of the neuron and $d\in\mathbb{N}^{+}$ is the dilation rate. The channel-wise PSN has identical FLOPs to the sliding PSN, and the latter can be regarded as a simplified case with setting $W[0][i]=W[1][i]=...=W[C-1][i]$ and $d=1$ in Eq.(10).
Additionally, when using multiple layers of dilated convolutions, setting the same dilation rate will lead to the grid effect. An approach to solving this issue is to assign the dilation rate using a sawtooth wave-like heuristic (Wang et al., 2018). Specifically, when constructing SNNs with channel-wise PSNs, we start by initializing with $d^{0}=1$ for the first spiking neuron layer, where the superscript represents the spiking neuron layer index. Then we update $d^{l}$ as:
@@
This approach ensures that after stacking multiple layers, the convolution in the time domain could effectively incorporates inputs from all time-steps. Figure 2 shows how the temporal receptive field increases with depths with (a) fixed dilations $d=1$ in the sliding PSN and (b) sawtooth dilations in the channel-wise PSN. The order is $k=2$ in both cases. It can be found that, with increasing depths, both neurons achieve larger temporal receptive fields. However, the sliding PSN can only capture the last 4 time-steps with 4 layers, while the channel-wise PSN can capture the last 7 time-steps.

### v2

# Multiplication-Free Parallelizable Spiking Neurons with Efficient Spatio-Temporal Dynamics

@@ Abstract
Spiking Neural Networks (SNNs) are distinguished from Artificial Neural Networks (ANNs) for their complex neuronal dynamics and sparse binary activations (spikes) inspired by the biological neural system. Traditional neuron models use iterative step-by-step dynamics, resulting in serial computation and slow training speed of SNNs. Recently, parallelizable spiking neuron models have been proposed to fully utilize the massive parallel computing ability of graphics processing units to accelerate the training of SNNs. However, existing parallelizable spiking neuron models involve dense floating operations and can only achieve high long-term dependencies learning ability with a large order at the cost of huge computational and memory costs. To solve the dilemma of performance and costs, we propose the mul-free channel-wise Parallel Spiking Neuron, which is hardware-friendly and suitable for SNNs’ resource-restricted application scenarios. The proposed neuron imports the channel-wise convolution to enhance the learning ability, induces the sawtooth dilations to reduce the neuron order, and employs the bit-shift operation to avoid multiplications. The algorithm for the design and implementation of acceleration methods is discussed extensively. Our methods are validated in neuromorphic Spiking Heidelberg Digits voices, sequential CIFAR images, and neuromorphic DVS-Lip vision datasets, achieving superior performance over SOTA spiking neurons. Training speed results demonstrate the effectiveness of our acceleration methods, providing a practical reference for future research. Our code is available at Github.
@@ 1 Introduction
Inspired by the biological neural system, Spiking Neural Networks (SNNs) are regarded as the third-generation neural network models [1]. By emulating neuronal dynamics and spike-based communication characteristics in the brain [2], SNNs effectively capture temporal information and achieve event-driven efficient computation, providing a novel paradigm for building the spike-based machine intelligence [3].
@@
Spiking neurons are the key component that distinguishes SNNs from Artificial Neural Networks (ANNs) [4]. They integrate input currents from synapses to membrane potentials by complex neuronal dynamics and fire spikes when the membrane potentials cross the threshold. These proceedings are generally described by several discrete-time difference equations [5, 6] in a formulation similar to the Recurrent Neural Networks. The discrete threshold-triggered firing mechanism induces the nondifferentiable problem and restricts the application of gradient descent methods. Recently, this problem has been resolved to a considerable degree by the surrogate gradient methods [7, 8, 9].
@@
Deep SNNs [10, 11, 12] commonly use stateless synapses, i.e., the weights are shared across time-steps and outputs only depend on the inputs at the same time-step, to extract spatial features. Consequently, dynamic spiking neurons play a critical role in SNNs in extracting temporal features, and this has attracted many research interests. Most of the previous research focuses on increasing the neuron model complexity with learnable parameters [5, 13] or adaptive dynamics [14], which strengthens the model capacity but brings extra computation costs, making it unfriendly for resource-restricted neuromorphic hardware. Another issue is that traditional spiking neuron models run in a serial step-by-step mode, limiting the utilization of the powerful parallel computing capabilities of Graphics Processing Units (GPUs) and resulting in a slower training speed for SNNs compared to ANNs. Recently, parallelizable spiking neuron models [15, 16, 17, 18] have been proposed that overwhelm traditional serial models in running speed, showing a promising solution to accelerate the training of SNNs.
@@
One of the most attractive characteristics of SNNs is that the multiply-accumulate (MAC) operations between binary spikes and synaptic weights can be superseded by accumulate (AC) operations during inference in neuromorphic chips [19]. However, in previous designs of spiking neurons, the computational costs of the neuronal dynamics have not been paid much attention. For instance, the Complementary Leaky Integrated-and-Fire (CLIF) neuron [14] introduces the computationally expensive sigmoid exponentiation, the Parallel Spiking Neuron (PSN), and the masked PSN [15] rely on the dense floating-point matrix multiplication. Compared to PSN and masked PSN, sliding PSN [15] only requires convolutional operations and demonstrates superior performance in handling time series with variable lengths. However, according to the study by [15], sliding PSN only achieves comparable performance to PSN when using a large convolutional kernel size, denoted as the order of the neuron, which significantly increases computational cost and memory usage. Moreover, these neurons still rely on massive multiply-accumulate (MAC) operations between floating-point neuronal weights and input currents.
@@
$1)$ To enhance the temporal information-capturing ability, we derive the sliding PSN by applying the channel-wise separable convolutions. To solve the dilemma of large temporal receptive fields and computational costs, we use dilated convolution. Compared to sliding PSN, our improvement does not introduce any additional floating point operations (FLOPs) and significantly reduces the inference memory.
[…]
$3)$ We discuss the implementations of SNNs with mul-free channel-wise PSN on GPUs for efficient training. We propose an autoselect algorithm to choose the fastest implementations, which is practical for future research about accelerating parallelizable spiking neurons.
$4)$ We achieve superior performance over other SOTA spiking neurons on various temporal tasks, validating the superior capability of the proposed methods in learning long-term dependencies.
@@ 2 Related Work
Hardware-friendly Network Design.
@@
To deploy neural network models to edge devices such as mobile phones and Field Programmable Gate Arrays with limited energy, memory, and computational ability, a promising solution is hardware-friendly network design. Various methodologies have been proposed. Network quantization [21] quantizes the original weights and activations to low bits. Typical models in 8-bit integers require $4\times$ less memory consumption and achieve up to $4\times$ faster computation than models in 32-bit floating-point. Network pruning [22] prunes synapses and neurons to reduce the size of the model. Classic pruning methods can achieve $4\times$ compression ratio for ResNet-18 on ImageNet with about $5\%$ accuracy drop [23]. Knowledge distillation [24] employs a large teacher network to supervise the learning of a small student network, and the student network can obtain a competitive or even higher performance than the teacher network. Apart from the above universal methods, SNNs [1, 25] achieve extreme power efficiency by asynchronous event-driven computation in tailored neuromorphic chips. For instance, Intel Loihi [26] consumes $48\times$ energy efficiency than CPUs in solving the LASSO optimization problem; Tsinghua Tianjic [19] achieves up to $10^{4}$ times power efficiency over the Titan-Xp GPU when classifying the NMNIST dataset [27].
Spiking Neuron Models.
@@
The improvement of spiking neuron models provides a general method to upgrade SNNs, which attracts much interest in the research community. The Parametric Leaky Integrated-and-Fire (PLIF) spiking neuron [5] parameterizes the membrane time constant $\tau_{m}$ by a sigmoid function and can learn $\tau_{m}$ by gradient descent during training, showing better accuracy and lower latency than the traditional Leaky Integrated-and-Fire (LIF) neuron with fixed $\tau_{m}$. The Gated LIF (GLIF) neuron [13] assembles the learnable gate units to fuse different bio-features in the neuronal behaviors of membrane leakage, integration accumulation, and reset, achieving impressive performance by these rich neuronal patterns. The Complementary LIF (CLIF) neuron [28] introduces the complementary membrane potential into the LIF neuron, effectively capturing and maintaining information related to membrane potential decay. However, the sigmoid used in its neuronal dynamics cannot be removed during inference, which is costly for neuromorphic chips. The Parallel Spiking Neuron (PSN) family [15] and the Stochastic Parallelizable Spiking Neuron (Stochastic PSN) [16] are the first parallelizable spiking neuron models. These two models convert the iterative neuronal dynamics to a non-iterative formulation by removing the neuronal reset. Extending from PSN, several variants are proposed. The Parallel Multi-compartment Spiking Neuron (PMSN) [29] introduces multiple interacting substructures to enhance the learning ability over diverse timescales. The Parallel Spiking Unit (PSU) [30] adds a fully-connected layer with sigmoid activations inside the neuron to approximate the neuronal reset. These methods obtain performance gains over PSN in certain datasets, but increase the complexity of neuron models and slow down training speeds.
@@ 3 Preliminary
3.1 Traditional Spiking Neurons
In general, spiking neurons can be described by three discrete-time difference equations [5]:
@@ 3.1 Traditional Spiking Neurons
$V[t]$ $=\begin{cases}H[t]\cdot(1-S[t])+V_{reset}\cdot S[t],\text{hard reset}\\ H[t]-V_{th}\cdot S[t],\text{soft reset}\end{cases}.$ (3)
Eq.(1) illustrates the neuronal charging process, where $X[t]$ is the input current at time-step $t$, extracted from the original spike input via a synapse layer (e.g., convolutional neural network or multi-layer perceptron). $H[t]$ and $V[t]$ are the membrane potentials before charging and after resetting at time-step $t$, and $f$ is the charging equation specified for different spiking neurons. In Eq.(2), $\Theta(x)$ is the Heaviside step function, defined as $\Theta(x)=1$ for $x\geq 0$ and $\Theta(x)=0$ for $x<0$. When $H[t]$ exceeds the threshold $V_{th}$, spiking neurons will fire spikes, and the membrane potential is reset as in Eq.(3). There are mainly two types of reset methods: hard reset will force the $V[t]$ to $V_{reset}$, while soft reset will subtract $V_{th}$ from $V[t]$.
@@ 3.2 Parallel Spiking Neuron
Fang et al. [15] found that for commonly used spiking neurons with a linear sub-threshold dynamic Eq.(1), such as the Integrate-and-Fire (IF) neuron and the LIF neuron, the neuronal dynamics could be expressed in a non-iterative form after removing the reset equation Eq.(3):
@@
Fang et al. [15] extended Eq.(4) by setting $W[t][i]$ as a learnable parameter, and proposed the PSN with the following neuronal dynamics:
@@
where $T$ is the sequence length. For simplicity, we ignore the batch dimension. The PSN does not involve iteration over time-steps. The core computation of the PSN is the matrix multiplication, which is highly optimized on GPUs and can be computed in parallel. Modified from the PSN, the sliding PSN is proposed by [15] with hidden states generating from the last $k$ inputs by a shared weight $\bm{W}\in\mathbb{R}^{k}$ across time-steps, whose neuronal dynamics are as follows:
@@ 4.1 Channel-wise and Dilated Convolution
In PSN and sliding PSN, the weights of the neurons are shared across all channels. However, the visualization of feature maps from an SNN conducted by [5] implies that the difference between channels is huge, e.g., one channel extracts the edges and another channel extracts the background at all time-steps (refer to Figure S4 in [5] for more details). This coarse design concept of sharing weights across channels may fail to capture the subtle disparity of features in channels. To solve this issue, we extend the weights to channel-wise.
To capture long-term dependency, the sliding PSN must use a large order $k$, resulting in a significant rise in computational cost and memory usage. We overcome this issue through the dilated convolution [31], where the convolution no longer processes consecutive inputs (,X[t2],X[t1],X[t])(...,X[t-2],X[t-1],X[t]), but instead (,X[t2d],X[td],X[t])(...,X[t-2d],X[t-d],X[t]), with $d>1$ as the dilation rate.
@@
where $\bm{W}\in\mathbb{R}^{C\times k}$ is the learnable weight, $k$ is the order of the neuron and $d\in\mathbb{N}^{+}$ is the dilation rate. The channel-wise PSN has identical FLOPs to the sliding PSN, and the latter can be regarded as a simplified case with setting W[0][i]=W[1][i]==W[C1][i]W[0][i]=W[1][i]=...=W[C-1][i] and $d=1$ in Eq.(10).
Additionally, when using multiple layers of dilated convolutions, setting the same dilation rate will lead to the grid effect. An approach to solving this issue is to assign the dilation rate using a sawtooth wave-like heuristic [32]. Specifically, when constructing SNNs with channel-wise PSNs, we start by initializing with $d^{0}=1$ for the first spiking neuron layer, where the superscript represents the spiking neuron layer index. Then we update $d^{l}$ as:
@@
This approach ensures that after stacking multiple layers, the convolution in the time domain could effectively incorporate inputs from all time-steps. Figure 2 shows how the temporal receptive field increases with depths with (a) fixed dilations $d=1$ in the sliding PSN and (b) sawtooth dilations in the channel-wise PSN. The order is $k=2$ in both cases. It can be found that, with increasing depths, both neurons achieve larger temporal receptive fields. However, the sliding PSN can only capture the last 4 time-steps with 3 layers, while the channel-wise PSN can capture the last 7 time-steps.